# Assessing the Efficacy of Invisible Watermarks in AI-Generated Medical Images

### v1

@@ 1 Introduction
To address these ethical considerations, researchers have explored adding identifiers, such as invisible watermarks, to synthetic images to help users distinguish between real and synthetic ones. These watermarking techniques are designed to be imperceptible to the human eye. SynID [5], used in Imagen [6], is a technique that effectively marks synthetic images with invisible watermarks, making them identifiable to users. In addition, instead of adding watermarks to synthesized images, researchers are also working on adding invisible water marks directly to generative models [7].
@@ 2.2 The Invisible Watermarking
In our study, we employed a traditional hybrid watermarking method that combines Discrete Wavelet Transform (DWT) and Discrete Cosine Transform (DCT). We chose to used this frequency based method because of its robustness compared to more recent deep learning-based watermarking methods [9, 10]. In our experiments, the word ”synthetic” was used as watermark information and was embedded into a a binary sequence.
@@ 4.2 Watermarking Ensures Uncompromised Synthetic Data Quality
Dataset
Fidelity ↑ Variety ↑ FID ↓ Privacy ↑
A2
0.345 50.949 0.266 /
+ WM
0.348 50.993 0.266 /
StyleGAN A1
0.378 166.445 15.585 0.472
+ WM
0.377 166.369 15.584 0.472
LDM A1
0.091 83.736 30.117 0.985
+ WM
0.091 83.725 30.115 0.985
StyleGAN A2
0.429 175.523 13.769 0.650
+ WM
0.427 175.662 13.775 0.650
LDM A2
0.102 84.531 27.484 0.983
+ WM
0.102 84.538 27.482 0.983
@@
4.3 Watermarking Has an Influence on Synthetic Data Uitlity

### v3

@@ 1 Introduction
To address these ethical considerations, researchers have explored adding identifiers, such as invisible watermarks [5], to synthetic images to help users distinguish between real and synthetic ones. These watermarking techniques are designed to be imperceptible to the human eye. SynthID [6], used in Imagen [7], is a technique that effectively marks synthetic images with invisible watermarks, making them identifiable to users. In addition, instead of adding watermarks to synthesized images, researchers are also working on adding invisible water marks directly to generative models [8].
@@ 2.2 The Invisible Watermarking
In our study, we employed a traditional hybrid watermarking method that combines Discrete Wavelet Transform (DWT) and Discrete Cosine Transform (DCT). We chose to used this frequency based method because of its robustness compared to more recent deep learning-based watermarking methods [10, 11]. In our experiments, the word ”synthetic” was used as watermark information and was embedded into a binary sequence.
@@ 4.2 Watermarking Ensures Uncompromised Synthetic Data Quality
Dataset Fidelity ↑ Variety ↓ FID ↓ Privacy ↑
A2 0.345 50.949 0.266 /
+ WM 0.348 50.993 0.266 /
StyleGAN A1 0.378 166.445 15.585 0.472
+ WM 0.377 166.369 15.584 0.472
LDM A1 0.091 83.736 30.117 0.985
+ WM 0.091 83.725 30.115 0.985
StyleGAN A2 0.429 175.523 13.769 0.650
+ WM 0.427 175.662 13.775 0.650
LDM A2 0.102 84.531 27.484 0.983
+ WM 0.102 84.538 27.482 0.983
@@
4.3 Watermarking Has an Influence on Synthetic Data Utility